OpenStudy (anonymous):

I need this so bad! what is the third degree polynomial function such that f(0)=-18 and whose zeros are 1,2,and3

OpenStudy (anonymous):

f(x)=a(x-1)(x-2)(x-3) f(0)=a(-6) so a=3 the answer: f(x)=3(x-1)(x-2)(x-3)

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

yw